OpenStudy (anonymous):

The length of a rectangle is 4 inches more than its width, and its perimeter is 34 inches. Find the length and width of the rectangle. Find the width of the rectangle. 6 in. 6.5 in. 7.5 in.

OpenStudy (anonymous):

So, w is your width. Your length would be w+4, then right? There are 2 sides w and 2 sides w+4 and all those added up total 34. Can you come up with the equation?

OpenStudy (anonymous):

yeah Becki

OpenStudy (anonymous):

Just checking back in to see if you got this worked out.

OpenStudy (texaschic101):

you can use the perimeter formula of a rectangle... Perimeter (P) = 2L (length) + 2W (width) Length is 4 more then the width L = W + 4 Perimeter = 34 now lets sub.. P = 2L + 2W 34 = 2(W + 4) + 2W -- distribute through the parenthesis 34 = 2W + 8 + 2W -- combine like terms 34 = 4W + 8 -- subtract 8 from both sides 34 - 8 = 4W 26 = 4W -- divide by 4 26/4 = W 6 1/2 = W -- (or 6.5) now sub 6.5 in for W L = W + 4 L = 6.5 + 4 L = 10.5 or 10 1/2 check.. P = 2L + 2W 34 = 2(10.5) + 2(6.5) 34 = 21 + 13 34 = 34 (correct) W = 6.5 L = 10.5